### v2

# Detecting Throat Cancer from Speech Signals using Machine Learning: A Scoping Literature Review

Abstract
Introduction: Cases of throat cancer are rising worldwide. With survival decreasing significantly at later stages, early detection is vital. Artificial intelligence (AI) and machine learning (ML) have the potential to detect throat cancer from patient speech, facilitating earlier diagnosis and reducing the burden on overstretched healthcare systems. However, no comprehensive review has explored the use of AI and ML for detecting throat cancer from speech. This review aims to fill this gap by evaluating how these technologies perform and identifying issues that need to be addressed in future research.
Materials and Methods: We conducted a scoping literature review across three databases: Scopus, Web of Science, and PubMed. We included articles that classified speech using machine learning and specified the inclusion of throat cancer patients in their data. Articles were categorized based on whether they performed binary or multi-class classification.
Results: We found 27 articles fitting our inclusion criteria, 12 performing binary classification, 13 performing multi-class classification, and two that do both binary and multiclass classification. The most common classification method used was neural networks, and the most frequently extracted feature was mel-spectrograms. We also documented pre-processing methods and classifier performance. We compared each article against the TRIPOD-AI checklist, which showed a significant lack of open science, with only one article sharing code and only three using open-access data.
Conclusion: Open-source code is essential for external validation and further development in this field. Our review indicates that no single method or specific feature consistently outperforms others in detecting throat cancer from speech. Future research should focus on standardizing methodologies and improving the reproducibility of results.
keywords:
Throat Cancer , Machine Learning , Artificial Intelligence , Speech , Vocal Pathologies
journal: arxiv
[…]
[first]organization=Faculty of Engineering and Physical Sciences University of Leeds,city=Leeds, country=UK
[…]
[second]organization=Ear, Nose and Throat Department Leeds Teaching Hospitals NHS Trust,city=Leeds, country=UK
1 Introduction
1.1 Background
The World Health Organisation predicts that cases of throat cancer (cancer of the larynx, nasopharynx, oropharynx, and hypopharynx) will rise by 15% by 2040 [1]. Early diagnosis is key to successful treatment, with Cancer Research UK stating that approximately 90% of patients will survive for five years or more after diagnosis with stage one laryngeal cancer compared to only 30% of patients diagnosed with stage four [2]. Late diagnosis also changes treatment options, with later-stage cancers requiring more aggressive and invasive treatment than early-stage cancers [3, 4]. Head and neck cancers, including throat cancer, are more common in individuals in low socioeconomic groups; as such, it’s important that diagnostic tools are developed to reduce the cost of diagnosis [5].
Identifying cancers is challenging and highly dependent on timely symptom recognition. Common early symptoms of throat cancer include changes in voice, pain or difficulty swallowing, and lumps in the neck [6].
Upon presentation with symptoms, patients undergo diagnostic techniques such as clinical history assessment, voice evaluation by clinicians, nasendoscopy, laryngoscopy, and biopsy [7]. Nasendoscopy, a standard outpatient procedure, uses a fibre-optic endoscope inserted through the nose to view the larynx and hypopharynx. If abnormalities are detected, biopsies can be taken under local anaesthesia in an outpatient setting or during a laryngoscopy under general anaesthesia for a comprehensive assessment of the larynx [7].
There are many benign voice disorders that have overlapping symptoms with throat cancer, making them difficult to differentiate without these invasive tests [8, 9]. Once cancer has been ruled out, speech assessment can be used to assist in the diagnosis of voice disorders and to measure the degree of abnormality. Two main protocols are commonly used: GRBAS and the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V). GRBAS is a reproducible clinical assessment of abnormal voices commonly used in clinical practice. Various aspects of the patient’s voice are scored on a scale of zero to three, where zero is normal, one is mild, two is moderate, and three is severe, across five different vocal elements: grade, roughness, breathiness, asthenia, and strain [10]. CAPE-V is an assessment tool which requires patients to perform three speech tasks. Six speech features are assessed for each task - overall severity, roughness, breathiness, strain, pitch, and loudness. Each feature is rated on a scale from 0 (normal) to 100 (severe) [11]. These protocols are subjective and require expert knowledge.
With healthcare systems being under immense pressure, it’s important that resources are correctly and most efficiently allocated. In the UK, only approximately 4.3% of patients referred on the urgent suspected cancer pathway for head and neck cancers have a cancer diagnosis; if tools are able to detect non-cancer patients for more efficient referral, some of the load on healthcare systems could be eased [12].
[…]
It has been suggested that by using ML or AI, throat cancer may be able to be detected from patient speech [16, 17, 18]. This technology may be able to detect throat cancer sooner and help prioritize those patients who are most at risk while reducing the need for invasive and expensive diagnostic procedures. In this work, we perform a scoping literature review to document and analyse current ML and AI methods for the detection of throat cancer from speech.
This work is structured as follows: Section 2 details the methods used to conduct this search including search terms, inclusion and exclusion criteria, research questions, study selection, and data extraction. Section 3 gives an overview of our results as well as focusing on each of the research questions defined in Section 2. We conclude our work in Section 4.
1.2 Related Work
Current work in the area of using patient speech for pathology detection using AI generally focuses on neurological conditions, most commonly dementia and Parkinson’s disease [19, 20]. There have been several conference challenges focused on the detection of dementia from speech [21, 22, 23]. In these challenges, participants were tasked with classifying patients with Alzheimer’s dementia from non-Alzheimer’s patients. These challenges use a mixture of acoustic and linguistic features in classification. This area differs significantly from the detection of throat cancer, as neurological diseases often lead to a decline in linguistic ability [24]. However, since throat cancer is a structural disorder rather than neurological, there is no effect on a patient’s linguistic ability but only on the acoustic features of speech.
1.3 Objective
The objective of this literature review is to explore current work using AI and ML for detecting throat cancer from speech. We formulated the following research questions (RQ):
  • 1.
    RQ1: What machine learning and artificial intelligence methods have been used for the detection of vocal pathologies, including throat cancer, from patient speech?
  • 2.
    RQ2: What features of speech can be used to identify pathological speech, including throat cancer?
  • 3.
    RQ3: What are the strengths of the existing research, and what issues need to be addressed in future work?
2 Methods
2.1 Search Strategy
We derived a set of search terms from the above research questions; the search terms can be seen in Table 1. These terms were discussed and agreed upon with clinical and technical experts.
The literature search was conducted across three databases: Scopus, Web of Science, and PubMed and was run on 08/07/2024. The publication’s title, abstract, and keywords were searched using the terms shown in Table 1. These search terms were chosen to try to cover all of the topics that should be included in articles relevant to this literature search. These topics are cancer, throat, speech, detection, and machine learning. The terms used for cancer, throat, and speech were discussed with an experienced ENT consultant and the terms used for detection and machine learning were discussed with a technical expert to ensure that all relevant terms were included.
Cancer Throat Speech Detection Machine Learning
cancer* throat audio classifi* “machine learning”
carcinoma* larynx speech diagnos* “artificial intelligence”
tumour* laryngeal sound detection “neural networks”
tumor* “voice box” spectrograms ml
neoplasm* glottis voice ai
malignan* glottic vocal
“vocal cord” prosody
supraglottis acoustic
supraglottic
Table 1: The search terms used in the literature review. Terms within “ ” are found exactly not separated, a * in a term can be replaced by any characters (for example, classifi* can be classification, classifier, classifiers etc).
2.2 Inclusion and Exclusion Criteria
In this review we include research study designs and exclude review articles, case studies, and clinical trials. Articles were also excluded if they were not written in English and if a full version of the manuscript was not available. No restrictions were placed on publication dates.
To be included in this study, the articles had to adhere to the following inclusion criteria.
  • 1.
    Specifically state that patients with throat cancer are included in their dataset.
  • 2.
    Use machine learning or artificial intelligence methods to classify patients with a vocal pathology.
  • 3.
    Use speech recordings (or features obtained from speech) as the primary input to the classification system.
2.3 Study Selection
Using the search terms stated in Table 1 across Scopus, Web of Science, and PubMed, a total of 273 articles were found, and 10 additional articles were also identified as being a part of a relevant challenge. We imported all of the articles into Zotero and then used Rayyan to facilitate the removal of duplicates as well as the manual removal of an additional duplicate (n=106) [25, 26]. We screened all article titles and abstracts and excluded ineligible articles (n=138). We then conducted a full article screening and removed any ineligible or inaccessible articles (n=13), resulting in 27 articles being included. Of the articles excluded at full-text screening, the majority (n=9) were excluded because they did not explicitly state the inclusion of cancer patients in their dataset, often looking to precancerous conditions instead. Three articles had the wrong outcome, not classifying pathological speech, and one article was unavailable. Figure 1 shows a flowchart based on the PRISMA framework depicting the process used to identify and screen the articles found in the search [27]. The screening process was carried out by a single researcher.
Figure 1: The Prisma diagram shows the steps taken to obtain relevant articles for this literature search.
2.4 Data Extraction
Work in this area follows a typical workflow shown in Figure 2. Preprocessing steps are any processing on the raw signal prior to feature extraction or classification. Feature extraction is commonly performed prior to classification and involves manipulating the signal in some way to extract features; these features are often meaningful and reduce the input size into the classification system, although some works may classify the signal without extracting any features.
Figure 2: The typical pipeline used in classifying pathological speech using machine learning. Both preprocessing and feature extraction are optional steps that are not performed by all articles.
A single reviewer extracted the following data from each article:
  • 1.
    The source and full reference.
  • 2.
    Any preprocessing steps performed.
  • 3.
    The features used in classification.
  • 4.
    The classification methods used.
  • 5.
    The results obtained.
The results of all attempted models were extracted from each article. The most common metrics across all articles were accuracy, specificity, and sensitivity. We show the equations for these metrics below:
Accuracy=Correctly classified speechTotal speech $\text{Accuracy}=\frac{\text{Correctly classified speech}}{\text{Total speech samples}}$ (1)
Specificity=Correctly classified healthy speechTotal healthy $\text{Specificity}=\frac{\text{Correctly classified healthy speech}}{\text{Total healthy speech}}$ (2)
Sensitivity=Correctly classified pathological speechTotal pathological $\text{Sensitivity}=\frac{\text{Correctly classified pathological speech}}{\text{Total pathological speech}}$ (3)
Another metric often used in the papers using multi-class classification was unweighted average recall (UAR). This metric is calculated by averaging the recall value for each of the specific pathologies included in the dataset. Equation 5 shows how it is calculated, where N is the number of pathologies in the dataset and $R_{i}$ is the recall (Equation 4) of the ith pathology in the dataset.
Recall=Correctly classified speech of pathology iTotal number of pathology i $\text{Recall}=\frac{\text{Correctly classified speech of pathology i}}{\text{Total number of pathology i speech}}$ (4)
$\text{UAR}=\frac{\sum_{i=1}^{N}{R_{i}}}{N}$ (5)
3 Results and Discussion
3.1 Overview
In this section, we provide a detailed overview of the articles included in this study. We grouped the articles based on whether they performed binary or multi-class classification. The extracted data from each article can be seen in Tables 2 and 3 for articles performing binary and multi-class classification, respectively. In both tables, we give the best performance reported in each article in terms of accuracy (Equation 1), and the corresponding specificity (Equation 2), and sensitivity (Equation 3), as those were the most commonly reported metrics. We also report the UAR (Equation 5) for the multi-classification articles, as this was most commonly reported in these articles. We note that two papers performed both binary and multi-class classification and are therefore included in both tables [28, 29]. Of the remaining articles, 12 performed binary classification, and 13 performed multi-class classification. Additionally, one of the articles reviewed is our previous work; despite this, we aim to maintain objectivity in this review [30].
Paper Preprocessing Features Classification Method(s) Accuracy Sensitivity Specificity
Gavidia-Ceballos and Hansen [31] Wiener filter Spectral features
Acoustic features
HMM 90.75 X X
Ritchings et al. [32] X Fast Fourier transform MLP X X X
Godino-Llorente and Gomez-Vilda [33] Antialiasing filter
Windowing
Endpoint detection
MFCC MLP
LVQ
96 94.90 95.89
Ben Aicha and Ezzine [34] X Glottal Parameters ANN 96.9 X X
Ezzine et al. [35] X Glottal Parameters MLP 98 X X
Fang et al. [36] X MFCC DNN
GMM
SVM
99.14 X X
Kim et al. [16] Normalization Acoustic features
MFCC
STFT
Raw signals
SVM
XGBoost
LightGBM
ANN
CNN (1D and 2D)
85.2 78.0 93.3
Kwon et al. [18] X MFCC CNN
DT
87.88 94.12 81.25
Wang et al. [17] X MFCC DNN 86.11 77.78 88.89
Chen et al. [29] X MFCC
Chroma vectors
SVM 85.00 X X
Paterson et al. [30] Band-pass filter
Wiener filter
Wavelet filter
Endpoint detection
Normalization
Acoustic features SVM
DT
RF
LR
81.2 87.2 X
Za’im et al. [37] X MFCC OSLEM
NB
SVM
DT
92 100 58
Kim et al. [28] X MFCC SVM
LightGBM
ANN
96.51 93.94 X
Wang et al. [38] X Raw signals
Medical records
CNN 81.5 78.3 81.6
Table 2: A summary of the papers using binary classification. MFCC - Mel-frequency cepstral coefficients, STFT - short-term Fourier transform, SVM - support vector machine, XGBoost - extreme gradient boosting, LightGBM - light gradient-boosting machine, ANN - artificial neural network, CNN - convolutional neural network, DT - decision tree, RF - random forest, DNN - deep neural network, MLP - multi-layered perceptron, LVQ - learning vector quantization, GMM - Gaussian mixture models, HMM - hidden Markov model, OSLEM - Online Sequential Extreme Learning Machine, NB - Naive Bayes, LR - logistic regression
Paper Preprocessing Features Classification Method(s) Accuracy Sensitivity Specificity UAR
Verikas et al. [39] X Medical records
Acoustic features
Cosine transform coefficients
SVM 80.47 X X X
Arias-Londoño et al. [40] Voice activity detection
Normalization
Perturbation features
MFCC
Spectrum features
Complexity features
GMM
GBT
SVM
k-nn
RF
X 92 54 61
Bhat and Kopparapu [41] X Acoustic features
MFCC
Spectral features
BayesNet
RF
X 96.6 66 68.67
Chuang et al. [42] X MFCC DNN X 93.1 46 62.87
Degila et al. [43] X MFCC SVM
RF
NN
LR
X 89.4 54.0 71.30
Grzywalski et al. [44] Low-pass filtering
Cropping
Data augmentation
MFCC
Acoustic features
DNN X 89.4 66.0 71.2
Islam et al. [45] Segmentation Prosodic features
MFCC
Acoustic features
SVM X 94.9 20 59.77
Ju et al. [46] Endpoint detection Acoustic features
MFCC
Spectral features
MIL
SVM
Label propagation
Transductive SVM
X X X 60.67
Pishgar et al. [47] X MFCC SVM X 88.60 78.23 59.00
Pham et al. [48] Endpoint detection MFCC SVM
RF
k-nn
Gradient Boosting
68.48 X X X
Ramalingam et al. [49] X Spectral features
MFCC
Time domain features
CNN
RNN
93 96 18 X
Fang et al. [50] X Medical records
MFCC
DNN
GMM
87.26 X X 81.59
Miliaresi et al. [51] Normalization
Segmentation
MFCC
Acoustic features
Medical records
NN 57 X X X
Chen et al. [29] X MFCC
Chroma vectors
SVM
CNN
50.66 X X X
Kim et al. [28] X MFCC SVM
LightGBM
ANN
82.61 72.94 X X
Table 3: A summary of the papers using multi-class classification. MFCC - Mel-frequency cepstral coefficients, DNN - deep neural network, NN - neural network, SVM - support vector machine, MIL - multi-instance learning, GMM - Gaussian mixture model, GBT - gradient boosting tree, k-nn - k nearest neighbour, RF - random forest, LR - linear regression, CNN - convolutional neural network
Our review covers articles published between 1996 and 2024, with no date restrictions applied. The distribution of publication years shows a peak of 10 articles published in 2018 due to a challenge that year and an increasing trend in the last three years (2022-2024) with seven articles published. The majority of these articles were conference papers (n=16) compared to journal articles (n=11). However, we do note that the majority of the conference articles (n=10) are from a single challenge.
3.1.1 Datasets
Gathering real-world datasets is hard due to restrictions related to data privacy and the infrequent routine collection of relevant speech data. When datasets are collected, they are often small, including fewer than 50 cancer patients.
Table 4 shows the size of the dataset used in each article splitting the datasets into the number of healthy, cancer, and non-cancer pathologies reported by each article. We found that some articles combine cancerous and precancerous conditions, where the number of cancerous and precancerous conditions is made clear we put the number of precancerous samples in the “Non-Cancer Pathology(s)” column. There is a significant outlier in the number of samples used in the articles as Ezzine et al. [35] reports over 6000 samples. We investigated this further and found that in this article, the open-source dataset Saarbruecken Voice Database is used [52]. This dataset contains over 2000 samples ($\sim$2700) and as such the samples referred to by Ezzine et al. cannot be directly from the Saarbruecken Voice Database. We speculate that the authors may have augmented their dataset by splitting audio samples into smaller segments, though this is not explicitly clarified in the article. This discrepancy means that it is impossible to reproduce their results even with the available data. The next largest dataset is fromKwon et al. [18], with 491 total patients. Figure 3 shows the percentage of patients that were healthy, had cancer and had non-cancer pathologies in each article. Gavidia-Ceballos and Hansen [31] has the highest percentage of cancer patients in their dataset (66.7%) with Godino-Llorente and Gomez-Vilda [33] having the lowest percentage (2.2%).
Article Healthy Cancer Non-Cancer Pathology(s)
Gavidia-Ceballos and Hansen [31] 10 20 0
Ritchings et al. [32] 20 20 0
Godino-Llorente and Gomez-Vilda [33] 53 3 79
Ben Aicha and Ezzine [34] $>$100 101 0
Ezzine et al. [35] 3009 3009 0
Fang et al. [36] 60 48 354
Kim et al. [16] 45 50 0
Kwon et al. [18] 33 176 282
Wang et al. [17] 0 43 129
Chen et al. [29] 0 40 398
Paterson et al. [30] 38 36 2
Za’im et al. [37] 252 37 93
Kim et al. [28] 155 30 178
Wang et al. [38] 0 60 1940
Verikas et al. [39] 25 22 193
Arias-Londoño et al. [40] 50 40 110
Bhat and Kopparapu [41] 50 40 110
Chuang et al. [42] 50 40 110
Degila et al. [43] 50 40 110
Grzywalski et al. [44] 50 40 110
Ju et al. [46] 50 40 110
Islam et al. [45] 50 40 110
Pishgar et al. [47] 50 40 110
Pham et al. [48] 50 40 110
Ramalingam et al. [49] 50 40 110
Fang et al. [50] 0 84 173
Miliaresi et al. [51] 200 50 0
Table 4: The reported samples in each article are split into healthy, cancer, and non-cancer pathologies.
Figure 3: The percentage of samples for healthy, cancer, and non-cancer pathology(s) in each article. [34] removed as exact counts were not provided.
3.1.2 Preprocessing
Preprocessing is a critical step in data preparation for machine learning, aimed at enhancing the quality of data and improving model performance. In this study, we examined the preprocessing methods reported in the reviewed articles. We found that the majority of articles did not report any preprocessing steps (n=17). Of the articles performing binary classification, only four used preprocessing, each employing different methods. A comparison of the performance of articles using preprocessing versus those not using it revealed no significant difference. For multi-class classification, six articles used preprocessing, while nine did not. Again, no significant difference in results was observed between those that used preprocessing techniques and those that did not.
The most common preprocessing technique identified was endpoint detection (including cropping and voice activity detection), followed by normalization and filtering. Some articles did not specify the reasons for their choice of preprocessing methods, indicating that these choices may have been made on an ad-hoc basis rather than for specific reasons supported by the literature. Although Kim et al. [16] and Paterson et al. [30] state that normalization reduces any effect that the distance between the microphone and the patient may have. Godino-Llorente and Gomez-Vilda [33] state that endpoint detection is used to avoid processing periods of silence, and Paterson et al. [30] states that filtering is used for the removal of background noise.
3.2 RQ1 - Classification Methods
In this section, we discuss Research Question 1: What machine learning and artificial intelligence methods have been used for the detection of vocal pathologies, including throat cancer, from patient speech? We present and evaluate the methods used in the examined articles and discuss their merits and drawbacks.
Many of the articles used multiple classification methods, with an average of 2.2 methods being used per article. The most common are neural networks (of varying types) and support vector machines (SVM).
Neural networks were originally modelled on the human brain. These networks consist of layers of nodes, one input layer, one or more hidden layers, and one output layer; each node in the network is connected to others with a certain weight and threshold [53]. There are many different types of neural networks, but they are typically large and complex. SVMs are smaller and less complex than neural networks; they work by separating classes using planes with varying functions; this technique is generally good on small but complex data [54].
Figure 4 shows the breakdown of the classification methods used in the articles. For clarity, methods used only once were grouped into the ’other’ category, and all types of neural networks were combined into a single category. Several papers also combined different algorithms into ensemble methods although this is not shown in Figure 4.
Figure 4: The classification methods used across all articles found in this review. The different neural network methods (ANN, BayesNet, CNN, DBN, DNN, LVQ, MLP, NN, RNN, OSELM) were grouped, and any method used only once was put into the other category for ease of visualisation. NN - neural network, SVM - support vector machine, RF - random forest, GMM - Gaussian mixture model, k-nn - k-nearest neighbour.
Figure 5 shows the performance of each of the classification models included in the found articles. This includes all results from articles that use multiple methods.
Figure 5(a) shows the accuracy achieved by each classification method used in articles for binary classification. In the context of binary classification, neural networks and SVMs demonstrated similar performance. LightGBM and XGBoost appeared to perform worse, but due to the limited number of studies using these methods, no definitive conclusion can be drawn.
For multi-class classification, all methods performed similarly in terms of UAR, with neural networks achieving the best results at the extreme. These results can be seen in Figure 5(b).
(a) The accuracy achieved in the binary classifiers using different classification methods
(b) The UAR achieved in the multi-class classifiers using different classification methods
Figure 5: The results of different classification methods from binary and multi-class tasks.
While these accuracy metrics provide valuable insights, it is important to note that the dataset used plays a crucial role in the performance of each technique. Therefore, comparing articles that use different datasets is challenging.
However, the IEEE FEMH Voice Data Challenge 2018 provides a unique opportunity for direct comparison, as participants were given the same dataset. Participants were provided with 200 samples to be used in training (40 neoplasm, 60 phonotrauma, 50 vocal palsy, and 50 healthy) and models were then tested on an undisclosed test set of 400 samples [41]. Figure 6 shows the UAR found using different methods in the challenge articles. The ten articles submitted to this challenge [44, 46, 45, 40, 48, 42, 41, 49, 47, 43] indicate that all methods perform similarly with overlapping UAR distributions. SVM classifiers exhibited a wider range of performance compared to ensemble methods and neural networks. These results suggest that no single classification method consistently outperforms the others.
Figure 6: The UAR found in the articles submitted to the IEEE FEMH Voice Data Challenge 2018 for different classification methods [44, 46, 45, 40, 48, 42, 41, 49, 47, 43].
3.3 RQ2 - Feature Extraction
In this section, we discuss Research Question 2: What features of speech can be used to identify pathological speech, including throat cancer?
We present the features extracted from the examined articles and discuss how they perform. In Figure 7, we show the number of each feature extraction method used in the articles. For clarity, methods used only once are grouped into the ’other’ category. Some articles combined features as input into classification systems although this is not shown in Figure 7.
Figure 7: The features used across all articles found in this review. Any feature used only once was put into the other category for ease of visualisation. MFCC - Mel-frequency cepstral coefficients.
The most commonly used feature is Mel-frequence cepstral coefficients (MFCC). MFCCs represent the spectrum of an audio signal and are often treated as images when being input into classification systems. MFCCs are calculated using three steps [33]:
  1. 1.
    Transforming the signal from the time to frequency domain using the short-time Fast Fourier Transform
  2. 2.
    Finding the energy spectrum of each frame and calculating the energy in each mel window
  3. 3.
    Taking logarithms and applying the cosine transform
The second most common feature extracted was acoustic features which are extracted directly from the audio signal without prior transformation. Common acoustic features include jitter, shimmer, and harmonic features [55, 56]. Articles that use these two features often source work in other domains as a reason for their use.
Figure 8 shows the results obtained from methods using different features. Figure 8(a) shows a disparity between the two most common feature sets in binary classification, with acoustic features performing consistently worse than MFCC. Glottal parameters seem to generally perform the best, although their performance still lies within the boundary of the articles using MFCC. As there is only one example of raw signals and STFT being used, it is hard to comment on their performance compared to the other methods.
Figure 8(b) shows the UAR achieved using different features in the multi-class tasks. These notably use a less wide range of features than the binary classification tasks, with a combination of features being used in many of the articles. Combining features seems to perform slightly better than MFCCs alone, although the performances are not significantly different.
As stated in the previous section, comparing the results of articles that use different datasets is hard. However, we can compare the results of the articles submitted to the IEEE FEMH Voice Data Challenge 2018. Figure 9 shows the results obtained based on the features used. In the challenge articles, only two types of features were used: MFCC and a combination of different features. This combination generally included MFCC, acoustic features, and spectral features. There is no clear advantage to either of these feature inputs.
From these results, we can see that no single feature set performs the best. For binary classification, glottal parameters have the best average accuracy, although this still lies within the range of results found when using MFCCs. The results of the IEEE FEMH Voice Data Challenge 2018 does not show any advantage when combining features over using MFCC alone.
(a) The accuracy achieved in the binary classifiers using different features
(b) The UAR achieved in the multi-class classifiers using different features
Figure 8: The results of different feature extraction methods from binary and multi-class tasks.
Figure 9: The UAR found in the articles submitted to the IEEE FEMH Voice Data Challenge 2018 for different input features [44, 46, 45, 40, 48, 42, 41, 49, 47, 43].
3.4 RQ3 - Strengths, Weaknesses and Recomendations
In this section, we address Research Question 3: What are the strengths of the existing research, and what issues need to be addressed in future work? To answer this, we compare each paper against the Transparent Reporting of a multivariable prediction model for Individual Prognosis Or Diagnosis (TRIPOD) AI checklist [57]. The TRIPOD-AI checklist sets out guidelines for the reporting of studies using AI or ML for the diagnosis or classification of medical conditions. We do acknowledge that this study was published in January 2024 and, as such, was not available to the authors of many of the studies found in this search. However, we still feel this is a good baseline against which the articles can be compared. TRIPOD-AI is split into 27 sections, each with several subsections totalling 50 points to be found in a study, although some of these points are not applicable to the studies in this search. Figure 10 shows the number of studies that included each of the points; for clarity, any points that were found to be not applicable to any of the studies have been removed. This figure shows the individual points that are defined in the TRIPOD-AI checklist, as well as the sections and subsections that they belong to. Table 5 shows the percentage of points each article achieved split by sections; we excluded sections that were not applicable for each article. We summarise our findings in this section split into key weaknesses and key strengths.
Figure 10: The results of comparing the articles to the TRIPOD-AI checklist. Any points that were not applicable to any of the articles in this study were removed. The labels on the left show the sections of the TRIPOD-AI checklist that each item was listed under, and the labels on the right show the subheading for each of the items.
Introduction Methods Open Science Results Discussion
Kim et al. [16] 50% 71% 1% 60% 40%
Wang et al. [17] 50% 82% 0% 60% 40%
Godino-Llorente and Gomez-Vilda [33] 25% 44% 0% 40% 0%
Ezzine et al. [35] 50% 29% 1% 20% 20%
Fang et al. [36] 50% 44% 0% 40% 60%
Ben Aicha and Ezzine [34] 25% 17% 0% 0% 20%
Ritchings et al. [32] 0% 12% 1% 0% 0%
Gavidia-Ceballos and Hansen [31] 25% 28% 0% 0% 20%
Kwon et al. [18] 100% 67% 1% 20% 60%
Grzywalski et al. [44] 50% 47% 0% 20% 0%
Miliaresi et al. [51] 25% 41% 1% 20% 40%
Fang et al. [50] 25% 50% 1% 60% 80%
Verikas et al. [39] 50% 39% 2% 40% 40%
Ju et al. [46] 50% 44% 0% 20% 40%
Islam et al. [45] 75% 44% 1% 20% 0%
Arias-Londoño et al. [40] 75% 39% 1% 20% 40%
Pham et al. [48] 75% 39% 2% 20% 40%
Chuang et al. [42] 50% 33% 2% 20% 20%
Bhat and Kopparapu [41] 25% 28% 0% 20% 0%
Ramalingam et al. [49] 0% 22% 2% 0% 20%
Pishgar et al. [47] 75% 33% 3% 20% 20%
Degila et al. [43] 75% 33% 0% 20% 60%
Wang et al. [38] 50% 50% 11% 20% 80%
Kim et al. [28] 75% 72% 9% 60% 60%
Chen et al. [29] 25% 56% 6% 60% 20%
Za’im et al. [37] 75% 72% 18% 60% 60%
Paterson et al. [30] 75% 44% 43% 60% 60%
Table 5: A table showing the percentage of points in the TRIPOD-AI checklist hit by each article of those applicable to that article.
3.4.1 Key Weaknesses
When comparing the papers to TRIPOD-AI we observed notable discrepancies between journal papers (average coverage of 18 points) and conference papers (average coverage of 11 points). We speculate that this disparity may arise from the typically stricter page limits imposed on conference papers compared to journal articles.
A critical area of weakness identified through TRIPOD-AI assessment is the lack of adherence to open science (points 18a-18f, and 19). Shockingly, only one of the rewieved articles has made its code available to the general public [30]; the lack of code sharing severely reduces the ability of other researchers to reproduce their results or validate their methods using external data. Furthermore, only three articles utilize open-source datasets [35, 38, 30]. Additionally, many articles fail to report conflicts of interest (point 20) and provide details on study protocol (point 22) or registration (point 22). While some articles mention data availability upon request, this practice is considered inadequate in promoting transparency.
Another significant issue is lack of information about the participants within the datasets, with point 6a being covered by six articles [16, 17, 18, 28, 29, 37], 6b being covered by five articles [16, 17, 18, 28, 29], and 6c being covered by six articles [16, 17, 18, 38, 28, 29]. These points cover a discussion of the study setting, eligibility criteria for participants, and details of any treatments received. We do acknowledge that this information is not always available when using data provided by a third party; however, we feel that this calls for better data descriptions to be provided when data is disseminated. We also believe that it is important to acknowledge the missing information about the data when discussing the limitations of the work. We have found that the articles, in general, are hard to compare due to the diversity in data sources and evaluation metrics. This is made worse by the lack of open data sharing, as subsequent studies cannot use the same data sources and be able to compare directly to previous studies.
Another area where papers fall short when compared to the TRIPOD-AI checklist is the lack of discussion on the practical application of AI systems in real-world medical settings (points 27a-27c). While half of the articles discuss steps for further research, only one addresses the impact of poor-quality input data, and none of the articles discuss the user inputs to a possible system. This gap in research indicates a need for more comprehensive studies that consider the practical implementation of AI systems in real-world medical settings, although this may be due to the relative newness of this area of study.
Additionally, the dependence on small datasets is concerning. As stated in Section 3.1, the largest reported dataset is that of Ezzine et al. [35], although we believe that this is an outlier due to omitted preprocessing steps, meaning that this large dataset does not represent the number of participants. The next largest dataset is that of Kwon et al. [18], containing 491 patients. Excluding Ezzine et al. [35] and Ben Aicha and Ezzine [34] due to unclear dataset size, the mean size of the datasets across the articles is 207 participants. We also note that only two of the included articles performs any external validation of their models [36, 38]; combining this with the small sample size puts them at high risk of bias and overfitting [58].
3.4.2 Key Strengths
On the other hand, strong areas identified through TRIPOD-AI analysis include robust reporting on data preparation and predictor definition (feature extraction techniques). Most papers include the steps for feature extraction or the equations used. We also find that the description of the healthcare context is generally strong, with clear motives for the work.
One of the key strengths in this area of study is the impressive results obtained by some of the articles. The binary classification accuracy ranges from 81-99%, showcasing the potential of the research. While the performance of the multi-class articles is harder to compare due to varying class and sample sizes, the variation in classification methods and feature extraction techniques demonstrates the promising future and potential for novel exploration in this field.
3.4.3 Recommendations
From these results, we can formulate some recommendations for future work in this area. It is clear that open science needs to be improved with future work, starting with code sharing; by providing a link to code availability within the article, other researchers are better able to understand all steps taken in the work and use it to improve their research. We believe that code sharing has very few disadvantages to authors. While we are aware of the limitations in making data available due to concerns around patient privacy and anonymity, we feel that better data descriptions and external validation on a publically available data source could allow for a better understanding of how these studies may be implemented in clinical practice and, alongside code sharing, allow for the validation of results. With the recent development of TRIPOD-AI, authors should attempt to adhere to the checklist where possible and acknowledge where they cannot provide information (for example, when information about a dataset is missing).
4 Conclusion
Within this scoping literature review, we identified 27 articles describing the classification of vocal pathologies using machine learning and speech recordings. Of these, 13 performed multi-class classification, 12 focused on binary classification, and two did both binary and multi-class classification. Our findings effectively addressed the research questions outlined in Section 1.3.
Neural networks emerged as the most commonly employed algorithm, although other algorithms demonstrated comparable success in classifying patients. The most common features used were MFCCs, although a wide variety of features were explored without a clear superiority observed among them. The results found in these works show that this is a promising area and that the use of AI in the detection of speech pathologies should be further investigated. We discuss the weaknesses in this area, particularly the lack of open science and the lack of publicly available code and datasets. The absence of publicly available code or datasets prevents the reproducibility and extension of findings across these articles. We also make recommendations for future research in this area to promote better open science and clarity of reporting. In future work, we will address this shortcoming, with the aim to create public code repositories, hence enabling external validation and facilitating improvements in subsequent studies.
Acknowledgements
This research was funded in part by the UKRI Engineering and Physical Sciences Research Council (EPSRC) [EP/S024336/1].